0.4n-18=3n-16-0.8h can you also show me how to check it?
Answer by checkley71(8403)

.4N-18=3N-16-.8N(?)
.4N-3N+.8N=-16+18
-1.8N=2
N=2/1.8
N=1/.9 OR 1 1/9 OR 1.11111 ANSWER.